OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 3x2 + 5x + 1 = 0

TheSmartOne (thesmartone):

\[\frac{ -5\pm \sqrt{25-4(3)(1)} }{ 2(1) }\]

OpenStudy (anonymous):

@TheSmartOne is correct good job

TheSmartOne (thesmartone):

\[\frac{ -5\pm \sqrt{25-12} }{2 }\] \[\frac{ -5\pm \sqrt{13} }{ 2}\]

OpenStudy (anonymous):

Thank you both.

OpenStudy (anonymous):

@TheSmartOne :wrong ..

TheSmartOne (thesmartone):

Oops I messed up

OpenStudy (anonymous):

Hmm?

OpenStudy (anonymous):

\[x= \frac{ -5\pm \sqrt{13} }{ 2(3) }\]

TheSmartOne (thesmartone):

Yes @Techworm I kind of lost the 3...

OpenStudy (anonymous):

I see now, thanks wormie worm.

OpenStudy (anonymous):

your welcome @Basher